OpenStudy (anonymous):

This is a three part question: f(x) = (5x+3)/(4x-9) g(x) = 3x/(4x-9) (f+g) (x) = ? domain of f+g = ? (f/g) (x) = ? domain of (f/g) = ?

OpenStudy (freethinker):

(f+g) (x) is the same as: [(5x+3)/(4x-9)] +[(3x/4x-9}] you have a common denominator so it is easy for you to perform the operation. try solving it from where I left off.

OpenStudy (anonymous):

(8x+3)/(4x-9)?

OpenStudy (jdoe0001):

yes

OpenStudy (freethinker):

ya! now what do you know about domains and ranges?

OpenStudy (anonymous):

that the x value can not produce 0 in the denominator?

OpenStudy (jdoe0001):

for a rational/fraction, the thing to understand is THAT, the denominator CANNOT be zero, \(\large \frac{0}{25}=0,\ but \frac{25}{0}=undefined \)

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so, the domain is those values for X, where X doesn't produce a zero in the denominator

OpenStudy (anonymous):

ok so...could this be all real numbers?

OpenStudy (jdoe0001):

well, one sec

OpenStudy (jdoe0001):

\(\large 4x-9=0 \implies x=\frac{9}{4} \)

OpenStudy (jdoe0001):

so, that "x" would give a 0, thus it cannot be part of the domain :)

OpenStudy (anonymous):

oh ok so could I say x =/ 9/4 or would I say x< 9/4 and x > 9/4

OpenStudy (anonymous):

and by =/ i mean \[x \neq 9/4\]

OpenStudy (jdoe0001):

\( (-\alpha, \frac{9}{4}) (\frac{9}{4}, +\alpha);\ x\ne \frac{9}{4} \)

OpenStudy (jdoe0001):

thereabouts

OpenStudy (anonymous):

you got it , that's a big gambit

OpenStudy (anonymous):

ok so how would i divide (5x+3) by 3x?

OpenStudy (jdoe0001):

ohh the second one, one sec

OpenStudy (jdoe0001):

well, tis simple really, when you have fractions as numerator and denominator, all you do is FLIP UPSIDE-DOWN the denominator fraction and multiply with the numerator :)

OpenStudy (jdoe0001):

multiplication is a simple 'straight across' multiplication

OpenStudy (anonymous):

so (5x+3)(4x-9) ?